QUESTION 3 Determine the total of the x-components of the three individual forces (F) 120 N 150 N 30° 80 N 40° 35°
Answer #1

The x component of the 80 N force which is applied at an angle of 40 with respect to x axis in the postive x-direction (toward right) is
F_{x,80}= F_{80}\cdot cos \ \theta = 80\cdot cos \ 40^{\circ}= 61.3 N
Similarly, The x-component of 120 N force which is applied at an angle of (40^{\circ}+30^{\circ})= 70^{\circ} with respect to x axis in the positive x-direction (rightward) is
F_{x,120}= F_{120}\cdot cos \ \theta = 120\cdot cos \ 70^{\circ}= 41 N
In the same way, The x-component of 150 N force which is applied at an angle of 35^{\circ} with respect to negative is
F_{x,150}= F_{150}\cdot cos \ \theta = 150\cdot cos \ (180^{\circ}-35^{\circ})=150\cdot cos\ 145^{\circ} =-122.8 N
The minus sign is for opposite direction with positive x axis or with the directions of F_{x,80} and F_{x,120} ..
Therefore the resultant force at the positive x direction is F_{x,80}+F_{x,120}+F_{x,150}=61.3+41-122.8=-20.57 N
The above value carries a minus sign so that the resultant force is 20.57 N at the negative x axis of toward left according to the diagram.